OpenStudy (anonymous):

If 30 N of force is exerted over an area of 20 m2, how much pressure is being applied?

OpenStudy (anonymous):

F = 20 N; S = 20 m2; p = F/S = 20 N/ 20 m2 p = 1 Pa